OpenStudy (anonymous):

how you simplify this improper fraction 10/5

OpenStudy (anonymous):

2 just divide

OpenStudy (anonymous):

\[\frac{ 10 }{ 5 } \div 5 =\frac{ 2 }{ 1 } = 2\]

OpenStudy (anonymous):

10/5 can divide evenly by 5 = 2/1 = 2